OpenStudy (anonymous):

Evaluate: 3m - 2x m = 5, x = 4

OpenStudy (anonymous):

\(\Large m = 5, x = 4 \\\Large \Rightarrow 3m-2x=3\times5-2\times4=15-8=7\)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Evaluate: -3r - 13 r = 2

OpenStudy (anonymous):

\(\Large -3\times2-13=-6-13=-19\)

OpenStudy (anonymous):

Evaluate: 3z + 2 z = 5

OpenStudy (anonymous):

\(\Large 3\times 5+2=15+2=17\)

OpenStudy (anonymous):

my friend was right lol was just making sure will yu help with my atha ones

OpenStudy (anonymous):

^^yes

OpenStudy (anonymous):

In this algebraic expression, what is the 3 called 6n + 3

OpenStudy (anonymous):

3 is called a constant term

OpenStudy (anonymous):

Evaluate: 5a + d - bc = a = 5, b = 7, c = 2, d = 4

OpenStudy (anonymous):

Evaluate: -3r - 13 r = 2

OpenStudy (anonymous):

1. \(\Large \color{red}{5a + d - bc =5\times5+4-7\times2=25+4-14=15}\) 2.\(\Large \color{red}{-3r - 13=-3\times2-13=-6-13=-19}\)

OpenStudy (anonymous):

thx i passed

OpenStudy (anonymous):

Glad to hear that^^

OpenStudy (anonymous):

=)